40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6973     Entry time: Fri Jul 13 13:02:52 2012
Author: Masha, Yaakov 
Type: Update 
Category: PEM 
Subject: GUR2 Fixed 

Yaakov and I investigated the GUR 2 problem. It turns out that the ADC channels that GUR 2 was plugged into, ADC channels 6 through 8 (on the actual ADC they are C7 through C9), did not correspond to the channels labelled "GUR 2" in the PEM, ADC channels 3 through 5. We modified them so that GUR 2 is now plugged into ADC channels 3 through 5 (on the ADC it's +1).

Before we discovered that this was the problem, we attempted to take the cover off of GUR 1 to check the gains, and discovered a stripped Allen screw on the side by the "Vertical" pot, which we removed.

Now the GUR 2 readout looks good, and we will give it more time to settle down before we take data.

 

ELOG V3.1.3-